# GEOcc: Geometrically Enhanced 3D Occupancy Network with Implicit-Explicit Depth Fusion and Contextual Self-Supervision

### v1

@@ Abstract
3D occupancy perception holds a pivotal role in recent vision-centric autonomous driving systems by converting surround-view images into integrated geometric and semantic representations within dense 3D grids. Nevertheless, current models still encounter two main challenges: modeling depth accurately in the 2D-3D view transformation stage, and overcoming the lack of generalizability issues due to sparse LiDAR supervision. To address these issues, this paper presents GEOcc, a Geometric-Enhanced Occupancy network tailored for vision-only surround-view perception. Our approach is three-fold: 1) Integration of explicit lift-based depth prediction and implicit projection-based transformers for depth modeling, enhancing the density and robustness of view transformation. 2) Utilization of mask-based encoder-decoder architecture for fine-grained semantic predictions; 3) Adoption of context-aware self-training loss functions in the pertaining stage to complement LiDAR supervision, involving the re-rendering of 2D depth maps from 3D occupancy features and leveraging image reconstruction loss to obtain denser depth supervision besides sparse LiDAR ground-truths. Our approach achieves State-Of-The-Art performance on the Occ3D-nuScenes dataset with the least image resolution needed and the most weightless image backbone compared with current models, marking an improvement of 3.3% due to our proposed contributions. Comprehensive experimentation also demonstrates the consistent superiority of our method over baselines and alternative approaches. The code will be made publicly available upon acceptance of this paper.
@@ I Introduction
3D vision for autonomous driving has undergone a significant shift from LiDAR-based multi-modal fusion to image-based vision-centric perception in recent years, primarily driven by the potential to reduce reliance on costly LiDAR sensors. The occupancy network has emerged as a cornerstone in vision-centric methodologies by transforming surround-view perspective images into a grid-based representation of the nearby environment. This includes both the top-down Bird-Eye-View (BEV) occupancy and voxel-like 3D occupancy formats. Compared to BEV occupancy, 3D occupancy enables a more fine-grained description of height information for various objects, e.g., poles, traffic signs, and bridges. Consequently, high-fidelity 3D representations depend on precise depth prediction and fine-grained semantic features from 2D images.
@@
For robust view transformation, explicit depth modeling (EDM) and implicit depth modeling (IDM) are separately adopted by mainstream occupancy networks. EDM employs a designated prediction network to extrapolate the depth probability distribution from pixels in images to a 3D space, represented by LSS [3], BEVDet [4] and OpenOccupancy [2]. The merit of EDM lies in its compartmentalized depth prediction process, facilitating precise depth estimation; however, the sparsity of LiDAR points limits the supervision of pixel depth. Conversely, IDM leverages cross-attention and self-attention mechanisms to directly transition image features to occupancy features, with implementations such as BEVFormer [5], TPVFormer [6], among others. IDM’s strength is in its enhanced differentiability and the generalizability of occupancy features, owing to its end-to-end transformation approach, though it can result in depth confusion. Thus, an integrated approach that combines explicit and implicit methods may offer a more effective solution by harnessing their respective advantages.
For sparse LiDAR supervision, the mainstream approach is to densify LiDAR points to cover more occupancy grids [2, 1] by mesh reconstruction, temporal accumulation, etc. However, inspired by current self-supervised depth estimation methods like VFDepth [7], we hypothesize that image-based self-supervision may alleviate sparse LiDAR issues in another novel approach. Given that occupancy features are derived from image features, self-supervision applied to the 2D perspective view can effectively propagate the supervisory signal back to occupancy features. Meanwhile, self-supervision can exploit information from multiple cameras and their temporal correlations, which largely extends the scope of LiDAR supervision signals.
@@ II Related Work
Vision-based 3D perception has been widely studied in recent years [8] due to its advantages in cost-effectiveness and straightforward data collection. Initial efforts focused on transforming 2D image features into Bird’s Eye View (BEV) planes for improved detection and planning capabilities, utilizing depth cues for precise 3D localization. The pioneering research LSS [3] and the following BEV-based detection methods [4, 5] explicitly predict the depth distribution of each pixel to project 2D features to 3D features. Subsequent research introduced implicit methods, evolving from BEV to full 3D occupancy perception [9, 6, 10, 11, 12, 13], employing attention mechanisms to learn view transformations. Despite these advancements, few have effectively integrated both implicit and explicit depth approaches for 3D occupancy under vision-only conditions. For example, FB-OCC [14] attempts to enhance explicit depth predictions using BEVFormer-based backward projection, yet its implicit BEV modeling omits vertical dimension details. Our work, GEOcc, aims to effectively merge the strengths of both depth paradigms. Unlike FB-OCC, GEOcc leverages a combined 3D occupancy representation through concerted implicit and explicit modules, complemented by an innovative occupancy compressor and a multi-scale Mask-based Encoder-Decoder structure to further refine occupancy features.
@@
The task of 3D occupancy prediction has garnered increasing attention in recent years, emerging from semantic scene reconstruction (SSC) tasks. Some previous work [15, 16, 17, 18]has explored the use of Occupancy grid maps to sense the environment for autonomous driving. However, due to the lack of semantic categories, simple 0-1 categorization does not provide a rich understanding of the environment. [19]introduced semantic segmentation and attempts to utilize raw radar data as input.[20] bring bayesian learning to occupancy tasks. 3D Occupancy Prediction area can be categorized into three primary research streams: LiDAR-based occupancy prediction (related to completion tasks) [21, 22, 23, 24], camera-only occupancy predictions [11, 14, 25, 26, 27, 12], and integrated LiDAR-Camera Occupancy predictions [2, 28]. LiDAR can serve as an informative prompt to substantially improve the accuracy of current occupancy prediction results enabling the forecasting of 4D occupancy dynamics over an imminent time frame [21, 23]. However, camera-only occupancy predictions are more challenging for accuracy improvement because of the robustness and generalizability bottleneck in 2D-to-3D view transformation without LiDAR input. COTR[29] focuses on compact representation, but lacks of rich 2D supervisory information on view transformations. To this end, our GEOcc focuses on vision-only occupancy predictions and attempts to geometrically enhance the view transformation process.
@@ III-B Network Structure
where $\mathbf{E}$ and $\mathbf{K}$ are camera extrinsic and intrinsic parameters and $\oplus$ means concatenation in feature dimension.
@@
where $\mathbf{X}_{l}\in\mathbb{R}^{N_{q}\times C}$ denotes $N_{q}$ semantic queries with $C$-dimension in $l$-th level and $\mathcal{M}_{l}$ denotes the binary mask predicted from previous layer. The linear transformations $f_{q}(\cdot)$, $f_{k}(\cdot)$, and $f_{v}(\cdot)$ remap semantic queries $\mathbf{X}_{l}$ and occupancy features $\mathbf{G}_{l}$ to new feature spaces as queries, keys, and values, respectively. Following the masked attention step, a Feed-Forward Network (FFN) layer is employed for feature projection. We use addition and normalization to construct a residual connection. Ultimately, the final query features $\mathbf{X}_{l}$ are processed by MLP to predict the semantic class $c_{i}$, and the corresponding binary 3D mask $\mathcal{M}_{i}$ for $c_{i}$ is generated by Eq. 4.
@@ III-C Self-supervised Geometric Pretraining
CTF-Occ [1] arXiv’2023 ResNet-101 928$\times$600 - 28.53
@@
PanoOcc [38] arXiv’2023 ResNet-101 640$\times$1600 - 42.13
OctreeOcc [27] arXiv’2023 ResNet-101 900$\times$1600 - 44.02
RenderOcc [12] arXiv’2023 Swin-B (88M) 512$\times$1408 - 26.11
@@ IV-A Experimental Setup
Dataset. Our experiments are conducted on the Occ3D-nuScenes [1] dataset, which comprises 700 scenes for training and 150 scenes for validation. Each frame contains LiDAR point cloud data of 32 lines and images captured by six surround-view cameras. The 3D occupancy ground-truths are provided at a resolution of $200\times 200\times 16$ voxels, encompassing a spatial range of [±$\left[\pm 40\text{m},\pm 40\text{m},-1-5.4\text{m}\right]$. Each voxel is a cube with an edge of 0.4m. The dataset delineates 17 semantic categories, with an additional $free$ label denoting unoccupied space.
@@
method Params(M) Latency(s)
BEVDet4D(R50) 35 0.41
Ours(R50) 64 0.45
FB-OCC(R50) 68 0.30
OccFormer(R50) 147 0.32
PanoOCC(R101) 98 0.55
BEVDet4D(SwinB) 122 1.42
Ours(SwinB) 151 1.46
@@
IV-B Main Results and the Discussion
In Tab. I, we present a comprehensive comparison of our method against leading methods on the Occ3d-nuScenes validation set. Furthermore, Tab. V shows a more detailed comparison for each semantic category
@@ IV-C Ablation Studies & Qualitative Visualization
Effectiveness of Each Component. The ablation study, presented in Tab. II, shows our baseline method achieving a 37.23% mIoU using EDM alone. Replacing EDM with IDM led to an approximate 10% drop in mIoU, while their combination yielded a 0.8% mIoU improvement, highlighting the advantage of integrating both depth modeling techniques. Incorporating the Mask-based transformer Encoder-Decoder (Masked-ED) structure further enhanced the model’s feature perception, resulting in a 1.80% mIoU increase. The implementation of Context-Aware Self-Training (CAST) loss improved performance by an additional 0 .25% mIoU, optimizing network initialization without increasing parameter count. Together, these enhancements contributed to a 3.27% mIoU improvement under the ResNet-50 backbone. Since the training time is very long, we use the smallest image resolution of 256x704 and the smallest image backbone ResNet-50 in the ablation experiments, and also remove the long-time sequence fusion module to reduce the training time. After the ablation experiments, we re-add the temporal fusion module, thus the final performance is consistent with that 43.64 reported in Tab. I.
@@
Figure 5: Visualization of occupancy prediction results under low light conditions and complex scenes. Better viewed when zoomed in.
@@
Ablation on Scales of Model Parameters. Table III shows the comparison between our method and other methods in terms of the number of model parameters and the amount of computation. For a fair comparison, all experiments were performed on a single NVIDIA RTX A6000 GPU, and the inference batch size is set to 1. The latency(s) is used to record computational time. Examining Tab. IV, a comparative analysis is conducted regarding the integration of our CAST into distinct training phases. It is evident that employing our CAST during the training phase significantly amplifies FLOPS, which inspires us to use a pretraining strategy.
@@
Qualitative Visualization Results Our method is compared with two high-performance benchmarks, BEVDet4D and FB-OCC, as illustrated in Figs. 4 and 6. In terms of geometric reconstruction, GEOcc provides superior continuity for slender and low-lying objects, such as tree trunks and fences, as shown by the green and gray grids, respectively, in Fig. 6. Semantically, BEVDet4D and FB-OCC both miss distant “sidewalks”—visible as dark purple bars in the BEV view in Fig. 4—whereas our method segments them accurately amidst large green “vegetation” areas, highlighting its semantic superiority. In Fig. 5 we visualize the occupancy prediction results in the lack of illumination and complex scenes, demonstrating the robustness of our method in specific scenes.

### v2

@@ Abstract
3D occupancy perception holds a pivotal role in recent vision-centric autonomous driving systems by converting surround-view images into integrated geometric and semantic representations within dense 3D grids. Nevertheless, current models still encounter two main challenges: modeling depth accurately in the 2D-3D view transformation stage, and overcoming the lack of generalizability issues due to sparse LiDAR supervision. To address these issues, this paper presents GEOcc, a Geometric-Enhanced Occupancy network tailored for vision-only surround-view perception. Our approach is three-fold: 1) Integration of explicit lift-based depth prediction and implicit projection-based transformers for depth modeling, enhancing the density and robustness of view transformation. 2) Utilization of mask-based encoder-decoder architecture for fine-grained semantic predictions; 3) Adoption of context-aware self-training loss functions in the pertaining stage to complement LiDAR supervision, involving the re-rendering of 2D depth maps from 3D occupancy features and leveraging image reconstruction loss to obtain denser depth supervision besides sparse LiDAR ground-truths. Our approach achieves State-of-the-Art performance on the Occ3D-nuScenes dataset with the least image resolution needed and the most weightless image backbone compared with current models, marking an improvement of 3.3% due to our proposed contributions. Comprehensive experimentation also demonstrates the consistent superiority of our method over baselines and alternative approaches. Our code is available at https://github.com/world-executed/GEOcc.git
@@ I Introduction
3D vision for autonomous driving [1, 2, 3, 4, 5, 6, 7] has undergone a significant shift from LiDAR-based multi-modal fusion to image-based vision-centric perception in recent years, primarily driven by the potential to reduce reliance on costly LiDAR sensors. The occupancy network has emerged as a cornerstone in vision-centric methodologies by transforming surround-view perspective images into a grid-based representation of the nearby environment. This includes both the top-down Bird-Eye-View (BEV) occupancy and voxel-like 3D occupancy formats. Compared to BEV occupancy, 3D occupancy enables a more fine-grained description of height information for various objects, e.g., poles, traffic signs, and bridges. Consequently, high-fidelity 3D representations depend on precise depth prediction and fine-grained semantic features from 2D images.
@@
For robust view transformation, explicit depth modeling (EDM) and implicit depth modeling (IDM) are separately adopted by mainstream occupancy networks. EDM employs a designated prediction network to extrapolate the depth probability distribution from pixels in images to a 3D space, represented by LSS [10], BEVDet [11] and OpenOccupancy [9]. The merit of EDM lies in its compartmentalized depth prediction process, facilitating precise depth estimation; however, the sparsity of LiDAR points limits the supervision of pixel depth. Conversely, IDM leverages cross-attention and self-attention mechanisms to directly transition image features to occupancy features, with implementations such as BEVFormer [12], TPVFormer [13], among others. IDM’s strength is in its enhanced differentiability and the generalizability of occupancy features, owing to its end-to-end transformation approach, though it can result in depth confusion. By combining both methods, we can not only address the distribution uncertainty in explicit depth estimation but also mitigate feature redundancy in implicit depth modeling.
For sparse LiDAR supervision, the mainstream approach is to densify LiDAR points to cover more occupancy grids [9, 8] by mesh reconstruction, temporal accumulation, etc. However, inspired by current self-supervised depth estimation methods like VFDepth [14], we hypothesize that image-based self-supervision may alleviate sparse LiDAR issues in another novel approach. Given that occupancy features are derived from image features, self-supervision applied to the 2D perspective view can effectively propagate the supervisory signal back to occupancy features. Through this self-supervised learning approach, the network can acquire rich geometric structural priors. Moreover, this prior information is denser than what is obtained through training directly on manually annotated ground truth. Meanwhile, self-supervision can exploit information from multiple cameras and their temporal correlations, which largely extends the scope of LiDAR supervision signals.
@@ II Related Work
Vision-based 3D perception has been widely studied in recent years [15] due to its advantages in cost-effectiveness and straightforward data collection. Initial efforts focused on transforming 2D image features into Bird’s Eye View (BEV) planes for improved detection and planning capabilities, utilizing depth cues for precise 3D localization. The pioneering research LSS [10] and the following BEV-based detection methods [11, 12] explicitly predict the depth distribution of each pixel to project 2D features to 3D features. These approaches often struggle with uncertainties in depth estimation, which can introduce errors during the feature transformation process. Additionally, the generated features are typically highly sparse, limiting their effectiveness in downstream networks. Subsequent research introduced implicit methods, evolving from BEV to full 3D occupancy perception [16, 13, 17, 18, 19, 20], employing attention mechanisms to learn view transformations. However, this methods may result in interactions between multiple voxels along the same ray and the same pixel, leading to feature redundancy and duplication. In contrast, our approach integrates explicit and implicit depth modeling and leverages 3D convolutions for effective feature fusion and compression. This combination significantly enhances feature robustness while mitigating redundancy. Despite these advancements, few have effectively integrated both implicit and explicit depth approaches for 3D occupancy under vision-only conditions. For example, FB-OCC [21] attempts to enhance explicit depth predictions using BEVFormer-based backward projection, yet its implicit BEV modeling omits vertical dimension details. Our work, GEOcc, aims to effectively merge the strengths of both depth paradigms. Unlike FB-OCC, GEOcc leverages a combined 3D occupancy representation through concerted implicit and explicit modules, complemented by an innovative occupancy compressor and a multi-scale Mask-based Encoder-Decoder structure to further refine occupancy features.
@@
The task of 3D occupancy prediction has garnered increasing attention in recent years, emerging from semantic scene reconstruction (SSC) tasks. Some previous work [22, 23, 24, 25]has explored the use of Occupancy grid maps to sense the environment for autonomous driving. However, due to the lack of semantic categories, simple 0-1 categorization does not provide a rich understanding of the environment. [26]introduced semantic segmentation and attempts to utilize raw radar data as input.[27] bring bayesian learning to occupancy tasks. 3D Occupancy Prediction area can be categorized into three primary research streams: LiDAR-based occupancy prediction (related to completion tasks) [28, 29, 30, 31], camera-only occupancy predictions [18, 21, 32, 33, 34, 19], and integrated LiDAR-Camera Occupancy predictions [9, 35]. LiDAR can serve as an informative prompt to substantially improve the accuracy of current occupancy prediction results enabling the forecasting of 4D occupancy dynamics over an imminent time frame [28, 30]. However, camera-only occupancy predictions are more challenging for accuracy improvement because of the robustness and generalizability bottleneck in 2D-to-3D view transformation without LiDAR input. COTR[36] focuses on compact representation, but lacks of rich 2D supervisory information on view transformations. To this end, our GEOcc focuses on vision-only occupancy predictions and attempts to geometrically enhance the view transformation process. Recent works, such as SparseOcc[37] and FlashOcc[38], have made progress in memory-efficiency. However, compared to GEOcc, they lack dense self-supervised pretraining and robust 3D feature representation. This is the key reason why GEOcc achieves a higher mIoU.
@@ III-B Network Structure
where $\mathbf{E}$ and $\mathbf{K}$ are camera extrinsic and intrinsic parameters and $\oplus$ means concatenation in feature dimension. Some traditional methods, such as BEVDet[11], BEVStereo[43], and OccFormer[18], rely solely on explicit depth estimation to lift 2D image features into 3D space. These approaches often suffer from uncertainties in depth estimation, leading to errors during the feature transformation process. Moreover, the resulting features tend to be highly sparse. Other methods, such as SurroundOcc[17] and BEVFormer[16], employ cross-attention between 3D and 2D for view transformation. However, this can result in interactions between multiple voxels along the same ray and the same pixel, causing feature redundancy and duplication. In contrast, our method combines both explicit and implicit depth modeling, and uses 3D convolutions for feature fusion and compression, thereby enhancing feature robustness.
@@
where $\mathbf{X}_{l}\in\mathbb{R}^{N_{q}\times C}$ denotes $N_{q}$ semantic queries with $C$-dimension in $l$-th level and $\mathcal{M}_{l}$ denotes the binary mask predicted from previous layer. The linear transformations $f_{q}(\cdot)$, $f_{k}(\cdot)$, and $f_{v}(\cdot)$ remap semantic queries $\mathbf{X}_{l}$ and occupancy features $\mathbf{G}_{l}$ to new feature spaces as queries, keys, and values, respectively. Following the masked attention step, a Feed-Forward Network (FFN) layer is employed for feature projection. We use addition and normalization to construct a residual connection. Ultimately, the final query features $\mathbf{X}_{l}$ are processed by MLP to predict the semantic class $c_{i}$, and the corresponding binary 3D mask $\mathcal{M}_{i}$ for $c_{i}$ is generated by Eq. 4. The mask-based semantic decoder demonstrates strong semantic segmentation capabilities. In works such as MaskFormer[45] and the subsequent Mask2Former[44], the semantic segmentation task is treated as a binary mask prediction for each category. Given its success in 2D image tasks, we introduce this approach to the 3D occupancy prediction task. We replace the original 2D attention mechanism with 3D attention and output 3D masks. Additionally, during the 3D feature encoding stage, we substitute the baseline 3D U-Net[11] structure with a Transformer-based model.
@@ III-C Self-supervised Geometric Pretraining
CTF-Occ [8] NIPS’2024 ResNet-101 928$\times$600 - 28.53
@@
PanoOcc [47] CVPR’2024 ResNet-101 640$\times$1600 - 42.13
OctreeOcc [34] CVPR’2024 ResNet-101 900$\times$1600 - 44.02
RenderOcc [19] ICRA’2024 Swin-B (88M) 512$\times$1408 - 26.11
@@ IV-A Experimental Setup
Dataset. Our experiments are conducted on the Occ3D-nuScenes [8] dataset, which comprises 700 scenes for training and 150 scenes for validation. Each frame contains LiDAR point cloud data of 32 lines and images captured by six surround-view cameras. Occ3D-nuScenes provides surround-view data captured by six cameras, making it highly suitable for our occupancy prediction task. Moreover, the dataset includes manually annotated occupancy ground truth, which facilitates supervised training. The 3D occupancy ground-truths are provided at a resolution of $200\times 200\times 16$ voxels, encompassing a spatial range of [±$\left[\pm 40\text{m},\pm 40\text{m},-1-5.4\text{m}\right]$. Each voxel is a cube with an edge of 0.4m. The dataset delineates 17 semantic categories, with an additional $free$ label denoting unoccupied space.
@@
IV-B Main Results and the Discussion
In Tab. I, we present a comprehensive comparison of our method against leading methods on the Occ3d-nuScenes validation set. Furthermore, Tab. V shows a more detailed comparison for each semantic category. The experiments in Tab. I and Tab. V involve the Long-Term Sequence Fusion Module and utilize two scales of image backbone: ResNet-50 and SwinTransformer-B.
@@ IV-B Main Results and the Discussion
method Params(M) Latency(s) mIoU(%)
BEVDet4D(R50) 35 0.41 40.43
Ours(R50) 64 0.45 43.64
FB-OCC(R50) 68 0.30 42.14
OccFormer(R50) 147 0.32 37.4
PanoOCC(R101) 98 0.55 42.13
BEVDet4D(SwinB) 122 1.42 43.29
Ours(SwinB) 151 1.46 44.67
@@ IV-C Ablation Studies & Qualitative Visualization
Effectiveness of Each Component. Since the training time is very long, we use the smallest image resolution of 256x704 and the smallest image backbone ResNet-50 in the ablation experiments, and also remove the long-term sequence fusion module to reduce the training time. The ablation study, presented in Tab. II, shows our baseline method achieving a 37.23% mIoU using EDM alone. Replacing EDM with IDM led to an approximate 10% drop in mIoU, while their combination yielded a 0.8% mIoU improvement. Our method combines explicit and implicit depth modeling, achieving superior feature robustness compared to these approaches. Additionally, for the classification head, our GEOcc employs a mask-based prediction head, enabling more precise predictions. Incorporating the Mask-based transformer Encoder-Decoder (Masked-ED) structure further enhanced the model’s feature perception, resulting in a 1.80% mIoU increase. The implementation of Context-Aware Self-Training (CAST) loss improved performance by an additional 0 .25% mIoU, optimizing network initialization without increasing parameter count. Our CAST module provides the network with rich geometric priors during the pretraining phase, enabling the model to gain an initial perception of spatial occupancy. This allows the model to achieve rapid convergence during the subsequent supervised training phase. Together, these enhancements contributed to a 3.27% mIoU improvement under the ResNet-50 backbone. After the ablation experiments, we re-add the temporal fusion module, thus the final performance is consistent with that 43.64 reported in Tab. I.
@@
Ablation on Scales of Model Parameters. Table III shows the comparison between our method and other methods in terms of the number of model parameters and the amount of computation. For a fair comparison, all experiments were performed on a single NVIDIA RTX A6000 GPU, and the inference batch size is set to 1. The latency(s) is used to record computational time.
We would like to point out that as all current Occupancy-based methods [12, 32, 36], which cannot achieve real-time performance, our method is not designed for online driving, although it shows great potential. However, the greater significance of GEOcc for autonomous driving lies in its ability to construct a high-precision offline ground truth system. Our method achieves the highest accuracy, making it ideal for simulation training and also service for autonomous driving.
Examining Tab. IV, a comparative analysis is conducted regarding the integration of our CAST into distinct training phases. It is evident that employing our CAST during the training phase significantly amplifies FLOPS, which inspires us to use a pretraining strategy.
Qualitative Visualization Results Our method is compared with two high-performance benchmarks, BEVDet4D and FB-OCC, as illustrated in Figs. 4 and 7. From the first row of Fig. 4, it can be seen that our method predicts more complete results for roadside trees, while the other methods show omissions. The second row demonstrates that our method accurately identifies roadside water barriers, whereas the other methods also exhibit missing predictions. These qualitative comparisons highlight the advantages of our method over the others. In terms of geometric reconstruction, GEOcc provides superior continuity for slender and low-lying objects, such as tree trunks and fences, as shown by the green and gray grids, respectively, in Fig. 4.
Visualization Analysis of extreme weather condition and low lighting scene We conduct further experimental analysis on the specific environmental conditions involved in the dataset. On the left of Fig. 5, raindrops adhering to the camera lens obscure imaging of vehicles and trees. On the right, rapid vehicle speed or road bumps result in motion blur, affecting the imaging of pedestrians along the roadside. Despite the disturbances caused by weather conditions and traffic situations, the visualization results show that our method achieves performance very close to the ground truth, with precise recognition of key objects. This demonstrates the robustness of our approach under challenging weather and traffic conditions. From the Fig. 6, we can observe that under low-light conditions, a camera positioned directly in front of the vehicle captures an image of a car located less than 40 meters ahead. The spatial occupancy grid predicted by our method clearly highlights the presence of the car, even in a nighttime setting with insufficient lighting and a noisy imaging environment.
Figure 5: Visualization analysis of model performance under low-quality imaging conditions. The left shows a rainy scene. The right illustrates motion blur caused by road bumps.
Figure 6: Visualization of occupancy prediction results under low light conditions.